OpenStudy (anonymous):

what are the next two terms 13, -5, -35, -77

OpenStudy (anonymous):

so the next two terms are -57 and -149

OpenStudy (anonymous):

thank you

OpenStudy (rsadhvika):

they cannot be next two terms

OpenStudy (anonymous):

?

OpenStudy (rsadhvika):

this looks tough, @amistre64 may be able to help if hez free :)

OpenStudy (rsadhvika):

wolfram saysm next two terms are -131, -197 but im not sure how to analytically get the answer http://www.wolframalpha.com/input/?i=+13%2C+-5%2C+-35%2C+-77%2C+...

OpenStudy (amistre64):

13, -5, -35, -77 -18 -30 -42 -12 -12 <-- this last row is a constant at -12

OpenStudy (amistre64):

these are known as difference tiers, each row represents the difference of the 2 elements above it if we can get down to a row thats constant, we can then develop a formula for any nth term

OpenStudy (anonymous):

so -12,-12 is the next two terms

OpenStudy (amistre64):

of course not ... but they do play a part in how we can construct the next 2 terms

OpenStudy (amistre64):

13, -5, -35, -77 a -18 -30 -42 b -12 -12 c c = -12 -42+c = b -77 + b = a

OpenStudy (anonymous):

-57,-149 ?

OpenStudy (amistre64):

-57 is not correct for the next term since it is obviously bigger than -77 and the terms are getting smaller

OpenStudy (anonymous):

but -149 is correct?

OpenStudy (amistre64):

13, -5, -35, -77 a -18 -30 -42 b -12 -12 c 13, -5, -35, -77 -77-42-12 -18 -30 -42 -42-12 -12 -12 -12

OpenStudy (amistre64):

13, -5, -35, -77 [-77-42-12] [-77-42-12-42-12-12] -18 -30 -42 -42-12 -42-12-12 -12 -12 -12 -12

OpenStudy (rsadhvika):

i get it, this is really cool method, thnks a lot @amistre64 :))

OpenStudy (amistre64):

-131 and -197 as has already been presented :)

OpenStudy (anonymous):

so its -131 and -197